OpenStudy (314dokg):

Hi guys. I just started learning Derivatives, and solving Derivatives with one variable is easy.. but with 2 is making me confuse... i got some sample questions here, and please try to explain how step by step. (i honestly don't know when to put in the dy/dx) Instruction: Find dy/dx x^3 + 2xy - y = 9 (y)/(xy - 1) = x y = (5)^(3x)^2

OpenStudy (luffingsails):

Can you not just solve the equations for x and then take the derivative? For instance... \[x^3 + 2xy-y=9\] \[x^3+y(2x-1)=9\] \[y=(9-x^3)/(2x-1)\] then perform your derivative...

OpenStudy (luffingsails):

Sorry, I meant solve for y.

OpenStudy (314dokg):

oh wait, forgot... the instruction was to find dy/dx... well my teacher made this process. which confuses me alot \[x^3y^2 + y =x\] \[x^3(2y)(\frac{ dy }{ dx })+y^2(3x^2)+1(\frac{ dy }{ dx }) = 1\] \[\frac{ dy }{ dx }(2x^3y+1)=-3x^2y^2\] \[\frac{ dy }{ dx } = \frac{ -3x^2y^2 }{ 2x^3y+1 }\]

OpenStudy (314dokg):

i just edited my question.... i must solve for dy/dx

OpenStudy (luffingsails):

So, the thing you have to remember is that y is a dependent variable (depended on x). Therefore, you treat all terms with y as having an x component. That's why why when you you did the first term of your example you solved it using the multiplication rule of derivatives.

ganeshie8 (ganeshie8):

Your teacher is using implicit differentiation, a very powerful technique.

OpenStudy (luffingsails):

\[x^3+2xy-y=9\] \[3x^2+2y+2x(dy/dx)=0\] \[dy/dx=(-2y-3x^2)/2x\] Seem right?

ganeshie8 (ganeshie8):

Your goal here is not to solve y, but to find the derivative of it, dy/dx. I said its powerful because it let's you find an expression for the derivative dy/dx without having to solve y first.

OpenStudy (314dokg):

so can you please explain briefly when is it the right time to place in the dy/dx if \[x^3 = 3x^2\] then what is y? or \[y^3\]

OpenStudy (314dokg):

and if they are together like \[x^4y^3\]

zepdrix (zepdrix):

You can always apply chain rule down to the variable you're differentiating with respect to. Example: \(\large\rm \frac{d}{dx}(x)^3=3(x)^2\frac{d}{dx}x\) I'm applying chain rule, multiplying by the derivative of the inner function. So we're left with: \(\large\rm =3x^2\frac{dx}{dx}\) But notice that dx/dx doesn't really have any significance. It's like asking, `how much does x change, as x changes?` So we always suppress this dx/dx when differentiating with respect to the variable we're dealing with. It doesn't give us any extra information. But I show you this just so you understand the pattern that the y's will follow.

zepdrix (zepdrix):

So when we apply this idea to something like this: \(\large\rm \frac{d}{dx}(y)^3=3(y)^2\frac{d}{dx}y\) Now our chain actually gives us something that we care about. \(\large\rm =3y^2\frac{dy}{dx}\) so we do not suppress it. And it all because we're differentiating with respect to some other variable.

zepdrix (zepdrix):

When you're given something like \(\large\rm (x^4y^3)'\) start by `setting up` your product rule, \(\large\rm =(x^4)'y^3+x^4(y^3)'\) And then deal with the derivatives one at a time. And again, since my ' is d/dx, we're only ending up with these derivative functions when we differentiate y (because that's the one that doesn't match our derivative operator). \(\large\rm =(4x^3)y^3+x^4(3y^2y')\)

OpenStudy (mww):

You'll get product or chain rule cases as zepdrix has explained. Independently though, In general \[\frac{ d }{ dx } f(y) = \frac{ d }{ dy } f(y) \frac{ dy }{ dx }= f'(y) \frac{ dy }{ dx }\]

OpenStudy (mww):

The equation I've written above illustrates why a dy/dx comes out as a factor when you differentiate an equation in y with respect to x.

OpenStudy (mww):

Here are a couple of examples just to give you an idea. The latter is just an application of product rule. \[\frac{ d }{ dx } (y^2) = \frac{ d }{ dy } (y^2) \frac{ dy }{ dx } = 2y \frac{ dy }{ dx }\] \[\frac{ d }{ dx }(3y) = \frac{ d }{ dy }(3y) \frac{ dy }{ dx } = 3\frac{ dy }{ dx }\] \[\frac{ d }{ dx } (xy) = x \frac{ d }{ dx }(y) + y \frac{ d }{ dx }(x) = x \frac{ dy }{ dx } + y\] or alternatively \[\frac{ d }{ dx }(xy) = \frac{ d }{ dy }(xy) \frac{ dy }{ dx } = (x \frac{ dy }{ dy } + y \frac{ dx }{ dy })\frac{ dy }{ dx }=x \frac{ dy }{ dx }+y\]